# A Survey on Large Language Models from General Purpose to Medical Applications: Datasets, Methodologies, and Evaluations

### v1

@@ III-E Summary
The choice of paradigm should be based on the demands, resources and time constraints of a specific project. In the medical field, where high expertise and accuracy are paramount, the paradigm CP-IFT-HA is more advantageous. However, in scenarios requiring rapid deployment with limited resources, the IFT paradigm can better meet practical needs. The CP-IFT and IFT-HA paradigms strike a balance between the performance of medical LLMs and the complexity of development. Specifically, the CP-IFT paradigm focuses on the comprehension and processing of medical knowledge, ensuring that the model can accurately grasp and apply relevant information. While the IFT-HA paradigm emphasizes the safety of responses and a doctor-like professionalism, ensuring patient safety and trust.
IV Evaluations
@@
Given the specialized and sensitive nature of medical information, erroneous or unsafe responses could potentially mislead patients, leading to harm. Hence, it is crucial to accurately and comprehensively assess the accuracy of medical LLMs. We systematically summarize the existing evaluation methods for medical LLMs and classify them into two main categories: Machine Evaluation and Human-Centric Evaluation. The former focus on the use of deterministic evaluation metrics to measure the performance of medical LLMs, while the latter evaluates them from a human perspective, considering subjective criteria. Tab. III presents the current evaluation protocols and metrics for medical LLMs.
@@
TABLE III: Evaluation Setting Details for the Medical Large Language Model. The abbreviations are as follows: STQA for Single-turn QA, MTQA for Multiple-turn QA, MCQA for Multiple-choice QA, NLP for Conventional Natural Language Processing Tasks , Zero. for Zero-shot Learning, Few. for Few-shot Learning, and SFT for Task-specific Supervised Fine-tuning.
@@
IV-A Machine Evaluation
IV-A1 Benchmarks
@@
Machine Evaluation assesses the performance of medical LLMs in handling natural language tasks within the medical domain, utilizing established benchmarks for natural language understanding and generation. These benchmarks provide explicit, static evaluation criteria that quantitatively measure the performance of medical LLMs.
@@
As shown in Tab. III, few studies have used natural language understanding tasks as evaluation benchmarks, such as Medical Named Entity Recognition, Medical Relationship Extraction, and Medical Text Categorization. General LLMs typically include natural language understanding tasks in training sets, providing them with capability in this area. This evaluation aims to determine whether Medical LLMs, after adaptation to medical information, retain natural language understanding ability in the medical domain. Me-LLaMA [50] and Taiyi [42] utilize natural language understanding benchmarks to evaluate the performance of the proposed medical LLM. They first perform task-specific supervised fine-tuning on the training sets of these benchmarks, followed by performance evaluation on the test sets. Common evaluation metrics for these tasks include Accuracy, F1-Score [128], and BERTScore [129].
@@
Many studies have used QA tasks (single-turn, multi-turn, multiple-choice) in natural language generation as evaluation benchmarks to measure the quality of response content and the performance processing medical QA tasks for medical LLMs. In particular, the single-turn QA benchmark involves a single question and a single answer, without context or continuous dialog. It allows direct assessment of the model’s performance and knowledge coverage on specific medical questions. And it is applicable to a wide range of medical topics, from simple definitions to complex diagnostic problems. The multi-turn QA benchmark involves consecutive questions and answers, such as a dialog between a doctor and a patient. This benchmark assesses the ability of medical LLMs to perform medical diagnosis, understand context (ensuring answers are based on previous dialog), and engage in interactive QA (testing the model’s ability to interact in real-world medical scenarios and maintain fluent dialog). The multiple-choice QA benchmark presents a question along with multiple options, requiring the model to identify the correct answer. With explicit correct answers, this benchmark allows precise quantification of the medical LLM’s capability. Additionally, some works not only evaluate the options outputted by the model but also assess reasoning content. The varied multiple-choice questions evaluate the model’s depth of knowledge across diverse medical domains and its logical reasoning ability. The commonly used evaluation metrics for natural language generation tasks are Accuracy, BLEU [130], ROUGE [131], Distinct [132] and GLEU [133]. In addition to traditional evaluation metrics, certain studies have introduced evaluation metrics tailored to healthcare dialog contexts. For instance, BianQue [39] proposes a novel evaluation metric named Proactive Questioning Ability (PAQ) to assess the capability of LLMs in proactively asking and querying questions to users.
@@
The majority of medical QA benchmarks comprise existing medical QA datasets, with only a limited number of benchmarks specifically gathered for particular evaluations. Note that the vast datasets employed in training LLMs may result in overlap between training and test sets. Therefore, data de-duplication and data leakage checks are necessary to ensure fairness of the test results.
IV-A2 Protocols
@@
The zero-shot protocol is to evaluate the ability of a model to directly perform a medical task with an unknown distribution, while assessing its generalization ability and adaptability. It only requires a natural language description of the downstream task, without providing specific examples, to ask the model to understand the medical task and use the intrinsic knowledge to accomplish this task. In the natural language understanding task, Me-LLaMA [50] and Taiyi [42] designed prompt templates for each downstream task to guide the output format of the model to meet the expected requirements. For the prompt for QA tasks, it usually give a role to the LLM and then guide the model’s output style in line with expectations. For instance, a prompt might be, ”You are an experienced doctor, please respond to the patient’s inquiries with patience and kindness. ### Patient: $<inputs>$## Doctor:”. Moreover, addressing medical issues often necessitates intricate multi-step reasoning. Hence, certain studies have employed prompt strategies to enhance model performance. For instance, MediTron [44] utilizes the prompt strategy Zero-shot Chain-of-Thought [134] to guide the model towards a more comprehensive understanding of medical problems, facilitating the generation of reasoning steps and answers. AntGLM-Med [46] introduces a simple prompt strategy called Verification-of-Choice (VoC), tailored for the multiple-choice QA benchmark. VoC operates under the assumption that each choice is correct, prompting the LLM to provide the corresponding explanation. Subsequently, these explanations serve as context for the LLM to identify the inconsistencies and give the final answer.
@@
Few-shot learning evaluates the performance of LLMs on downstream tasks involving a limited number of examples. Typically, examples are derived from the training set in the downstream task, including both inputs and corresponding outputs. The medical LLM then utilizes these examples, in conjunction with its intrinsic medical knowledge, to generate appropriate outputs in accordance with the provided instructions. Beyond simply employing a limited number of examples as context, several studies [44] have incorporated prompting strategies like Chain-of-Thought [135] and Self-Consistency Chain-of-Thought (SC-CoT) [136] to stimulate the reasoning capabilities of medical LLMs. These strategies aim to enhance the accuracy of outputs and the rationality of reasoning processes. Additionally, prompting strategies can facilitate the identification and correction of potential errors and knowledge gaps in medical LLMs, particularly in complex medical tasks.
@@
Task-specific supervised fine-tuning involves training a medical LLM on a benchmark training set and evaluating it on a corresponding test set [48, 41]. By training the medical LLM on these specific medical tasks, such as drug named entity recognition and medical literature comprehension, the generalization performance and knowledge transfer ability of the model can be evaluated. Two types of training approaches, parameter-efficient fine-tuning and full-parameter fine-tuning, are used to evaluate the performance of the model under two different levels of medical knowledge injections [49].
IV-B Human-Centric Evaluation
@@
Given the inherent diversity and complexity of model responses, machine evaluation methods may not adequately capture the performance of medical LLMs, particularly in assessing the usefulness and safety of outputs. To address this limitation, researchers have increasingly adopted Human-Centric Evaluation approaches. It provides a more realistic assessment of medical LLMs in real-world applications, reflecting human judgment and expectations.
IV-B1 Evaluator
@@
Existence studies employ both humans and LLMs as evaluators. Since the users of medical LLMs are humans, human evaluation can be effective and direct in assessing medical LLMs. Given the specialized nature of the medical field, existing work categorizes human evaluation into expert evaluation and lay user evaluation, reflecting the perspectives of physicians and general users, respectively. While human evaluation offers valuable insights into LLM performance, it is not without its limitations. Human evaluation can be time-consuming and resource-intensive, and the results can be susceptible to individual biases and expertise levels. To address these challenges, some researchers have explored employing advanced LLMs to replace human evaluation of medical LLMs. These advanced LLMs are trained on vast datasets and possess a knowledge base comparable to that of humans. Since the LLM is to evaluate the performance of medical LLMs from a human perspective, the LLM evaluation is included in Human-Centric Evaluation.
IV-B2 Evaluation Dimensions
@@
In Human-Centric Evaluation, the initial step involves identifying the evaluation dimensions for the model. Given the inconsistent terminology employed in previous studies of medical LLMs to describe these dimensions, we standardized the terminology and provided corresponding explanations. Tab. III presents the normalized terms.
@@
Doctor-like: Exhibits the communication style and tone of a professional physician, reflecting a comprehensive understanding of medical principles and clinical expertise. Capable of providing clear explanations of intricate medical concepts and offering well-founded diagnostic insights and recommendations.
@@
Patient-friendly: Interacts with users in a patient and empathetic manner, fostering a sense of understanding and support. Effectively simplifies complex medical concepts into easily comprehensible language, demonstrating compassion and care, thereby alleviating user anxiety and enhancing trust and satisfaction.
@@
Professional: Exhibits a high degree of professionalism, characterized by comprehensive medical knowledge and sound clinical judgment. Provides treatment recommendations and prescriptions that are not only accurate and reliable, but are also conform to medical norms and ethics ensuring the achievement of optimal patient outcomes.
@@
Safe: Responses are safe and harmless, and not geographically or racially discriminatory. Correctly address unhealthy and unsafe issues, provide accurate and reliable advice, and avoid misleading users or raising health risks. Always follow medical ethics, protect user privacy, and ensure the accuracy and safety of information delivery.
@@
Helpful: Capable of meeting the user’s needs for medical consultation or the physician’s clinical instructions to achieve the expectations of users.
Given the critical importance of safety in the medical field, this dimension must be assessed by human experts. Other dimensions can be evaluated by LLMs, thereby increasing efficiency. For non-professional users, the evaluation dimensions focus on Patient-friendly and Fluency. These two dimensions match the user’s perspective and can measure the real feedback of the model from users in real scenarios.
IV-B3 Protocols
Human-centric evaluation comprises two primary protocols: individual evaluation and pairwise evaluation. Individual evaluation involves a human evaluator or a LLM scoring the output based on predefined dimensions, with only a single question and its corresponding output visible. The scoring range is a predetermined interval, and the average score for each dimension is statistically calculated as the final result. This metric quantitatively reflects the model’s overall performance in each dimension, identifying strengths and weaknesses, and guiding future optimizations on weaker dimensions. Pairwise evaluation involves a human evaluator or a LLM comparing the outputs of two models based on predefined dimensions and selecting the better-performing one. Evaluation metrics include win rate (the proportion of evaluated models outperforming the comparison model), tie rate (the proportion of the two that are equivalent), and loss rate (the proportion of evaluated models underperforming the comparison model). In addition, to minimize the positional bias existing in the LLM as a judge, the positions of the two model outputs are exchanged for another comparison. Pairwise evaluation facilitates the rapid identification of the best-performing model from multiple candidates, serving as a foundation for model optimization and deployment.
@@ IV-B3 Protocols
V Challenges and Future Directions
@@
V-A Data Privacy
@@
To enable medical LLMs to achieve capabilities comparable to those of medical professionals, it is essential to integrate more real medical data into these models to enhance their medical knowledge. Data from professional medical organizations tend to be more authentic and reliable than data sourced from the web. However, due to the sensitive nature of medical data and associated policies, such data is seldom publicly available. Consequently, a significant challenge is to ensure the secure utilization of medical data within organizations while safeguarding patient privacy. One approach to addressing this issue is federated learning [137], a distributed training method that enables model parameters to be shared across multiple organizations without sharing the data itself. By keeping data stored locally, federated learning preserves data privacy while allowing the model to learn from diverse data sources, thus enhancing its generalization and performance. Federated learning has been applied to medical images [138], sensor data [139], and natural language text [140], promoting the secure use of health data among medical organizations.
@@
Currently, federated learning has several shortcomings in the medical field that require further investigation. Differential privacy [141] can be integrated with federated learning to enhance data privacy protection, but this integration degrades model performance. How to balance privacy protection and model performance is a concern. Additionally, the quality and quantity of data varies among healthcare organizations, and how to evaluate the contribution of each organization is a key topic. Finally, medical LLMs have a significantly higher number of parameters compared to pretrained language models, raising the issue of optimizing communication and computational overhead during the transmission and aggregation of parameters in large-scale models.
V-B Knowledge Augmentation

### v2

@@ 3.5 Summary
The choice of paradigms should be based on the demands, resources, and time constraints of a specific project. In the medical field, where high expertise and accuracy are paramount, the CP-IFT-HA paradigm is more advantageous. However, the IFT paradigm can better meet practical needs in scenarios requiring rapid deployment with limited resources. The CP-IFT and IFT-HA paradigms strike a balance between the performance of medical LLMs and the complexity of development. Specifically, the CP-IFT paradigm focuses on the comprehension and processing of medical knowledge, ensuring that the model can accurately grasp and apply relevant information. While the IFT-HA paradigm emphasizes the safety and doctor-like professionalism of responses, ensuring patient safety and trust.
4 Evaluations
@@
Given the specialized and sensitive nature of medical information, erroneous or unsafe responses could potentially mislead and harm patients. Hence, it is crucial to accurately and comprehensively evaluate the performance of medical LLMs [20, 127]. We systematically summarize the existing evaluation methods for medical LLMs and classify them into two main perspectives: Machine Evaluation and Human-Centric Evaluation. The former focuses on using deterministic evaluation metrics to measure the performance of medical LLMs, while the latter evaluates them from a human perspective. Tab. 3 presents the evaluation protocols and metrics for existing medical LLMs.
@@
Table 3: Evaluation Setting Details for the Medical Large Language Model. The abbreviations are as follows: STQA for Single-turn QA, MTQA for Multiple-turn QA, MCQA for Multiple-choice QA, NLP for Conventional Natural Language Processing Tasks, Zero. for Zero-shot Learning, Few. for Few-shot Learning, and SFT for Task-specific Supervised Fine-tuning.
@@
EpilepsyLLM [200] STQA BLEU, ROUGE, METEOR, SPICE Zero.
@@
Hippocrates [3] MCQA Accuracy Zero., Few.
MING-MOE [94] STQA, MCQA F1, ROUGE, Accuracy Zero. Human Case Study
Lingdan [67] STQA, MTQA Top@K Zero.,
Aloe [51] STQA, MCQA Accuracy, ASR Zero., Few.
PediatricsGPT [177] STQA BLEU, ROUGE, GLEU, Distinct Zero. Human, LLM Professional, Helpful, Safe, Fluent Pairwise, Case Study
Aqulia-Med [198] STQA, MTQA, MCQA Accuracy Zero., Few. LLM Professional, Fluent, Helpful Pairwise
4.1 Machine Evaluation
4.1.1 Benchmark Types
@@
Machine Evaluation assesses the performance of medical LLMs in handling natural language tasks, utilizing established benchmarks for natural language understanding and generation. These benchmarks provide explicit, static evaluation criteria for quantitatively measuring the performance of medical LLMs.
@@
As shown in Tab. 3, few studies have used natural language understanding tasks as evaluation benchmarks, such as Medical Named Entity Recognition, Medical Relationship Extraction, and Medical Text Classification. General LLMs typically include natural language understanding tasks in training sets, providing them with capability in this area. This evaluation aims to assess whether Medical LLMs retain natural language understanding ability after adaptation to medical information. Me-LLaMA [171] and Taiyi [102] utilize natural language understanding benchmarks to evaluate the performance of the proposed medical LLM. They first perform task-specific supervised fine-tuning on the training sets of these benchmarks, followed by performance evaluation on the test sets. Common evaluation metrics for these tasks include Accuracy, F1-Score [120], and BERTScore [192]. Additionally, a limited number of studies employed the evaluation metrics METEOR [10], SPICE [6], and Attack Success Rate (ASR) [51].
@@
Many studies have used QA tasks (single-turn, multi-turn, and multiple-choice) in natural language generation as evaluation benchmarks to measure the quality of response content and the performance of processing medical QA tasks for medical LLMs. In particular, the single-turn QA benchmark involves a single question and a single answer without continuous dialogue. It allows direct assessment of the model’s performance and knowledge coverage on specific medical questions. And it is applicable to a wide range of medical topics, from simple definitions to complex diagnostic problems. The multi-turn QA benchmark involves consecutive questions and answers, such as a dialogue between a doctor and a patient. This benchmark assesses the ability of medical LLMs to perform medical diagnosis, understand context (ensuring answers are based on previous dialogue), and engage in interactive QA (testing the model’s ability to interact in real-world medical scenarios and maintain fluent dialogue). The multiple-choice QA benchmark presents a question and multiple options, requiring the model to identify the correct answer. With explicit correct answers, this benchmark allows precise quantification of the medical LLM’s capability. Additionally, some works not only evaluate the options outputted by the model but also assess reasoning content. The varied multiple-choice questions evaluate the model’s depth of knowledge across diverse medical domains and its logical reasoning ability. The commonly used evaluation metrics for natural language generation tasks are Accuracy, BLEU [116], ROUGE [95], Distinct [86] and GLEU [169]. In addition to traditional evaluation metrics, certain studies have introduced evaluation metrics tailored to healthcare dialogue contexts. For instance, BianQue [23] proposed a novel evaluation metric named Proactive Questioning Ability (PAQ) to assess the capability of LLMs in proactively asking and querying questions to users.
@@
The majority of medical QA benchmarks comprise existing medical QA datasets, with only a limited number of benchmarks specifically gathered for particular evaluations. Note that the vast datasets employed in training LLMs may overlap between training and test sets. Therefore, data de-duplication and data leakage checks are necessary to ensure the fairness of the test results.
4.1.2 Protocols
@@
Zero-shot protocol is to evaluate the ability of a model to directly perform a medical task with an unknown distribution while assessing its generalization ability and adaptability. It only requires a natural language description of the downstream task, without providing specific examples, to ask the model to understand the medical task and use the intrinsic knowledge to accomplish this task. In the natural language understanding task, Me-LLaMA [171] and Taiyi [102] designed prompt templates for each downstream task to guide the output format of the model to meet the expected requirements. For the prompt of QA tasks, it usually gives a role to the LLM and then guide the model’s output style in line with expectations. For instance, a prompt might be, "You are an experienced doctor, and please respond to the patient’s inquiries with patience and kindness. ### Patient: $<inputs>$## Doctor:". Moreover, addressing medical issues often necessitates intricate multi-step reasoning. Hence, certain studies have employed prompt strategies to enhance model performance. For instance, MediTron [24] utilizes the prompt strategy Zero-shot Chain-of-Thought [78] to guide the model towards a more comprehensive understanding of medical problems, facilitating the generation of reasoning steps and answers. AntGLM-Med [89] introduces a simple prompt strategy called Verification-of-Choice (VoC), tailored for the multiple-choice QA benchmark. VoC operates under the assumption that each choice is correct, prompting the LLM to provide the corresponding explanation. Subsequently, these explanations serve as a context for the LLM to identify the inconsistencies and give the final answer.
@@
Few-shot protocol evaluates the performance of LLMs on downstream tasks involving a limited number of examples. Typically, examples are derived from the training set in the downstream task, including both inputs and corresponding outputs. The medical LLM then utilizes these examples, in conjunction with its intrinsic medical knowledge, to generate appropriate outputs in accordance with the provided instructions. Beyond simply employing a limited number of examples as context, several studies [24] have incorporated prompting strategies like Chain-of-Thought [162] and Self-Consistency Chain-of-Thought (SC-CoT) [158] to stimulate the reasoning capabilities of medical LLMs. These strategies aim to enhance the accuracy of outputs and the rationality of reasoning processes. Additionally, prompting strategies can facilitate the identification and correction of potential errors and knowledge gaps in medical LLMs, particularly in complex medical tasks.
@@
Task-specific supervised fine-tuning involves training a medical LLM on a benchmark training set and evaluating it on a corresponding test set [79, 180]. By training the medical LLM on these specific medical tasks, such as drug named entity recognition and medical literature comprehension, the generalization performance and knowledge transfer ability of the model can be evaluated. Two types of training approaches, parameter-efficient fine-tuning, and full-parameter fine-tuning, are used to evaluate the performance of the model under two different levels of medical knowledge injections [121].
4.2 Human-Centric Evaluation
@@
Given the inherent diversity and complexity of model responses, the machine evaluation may not adequately captures the performance of medical LLMs, particularly in assessing the usefulness and safety of outputs. To address this limitation, researchers have increasingly adopted Human-Centric Evaluation approaches. It provides a more realistic assessment of medical LLMs in real-world applications, reflecting human judgment and expectations.
4.2.1 Evaluator
@@
Existence studies employ both humans and LLMs as evaluators. Since the users of medical LLMs are humans, human evaluation can be effective and direct in assessing medical LLMs. Given the specialized nature of the medical field, existing work categorizes human evaluation into expert evaluation and lay user evaluation, reflecting the perspectives of physicians and general users, respectively. While human evaluation offers valuable insights into LLM performance, it is not without its limitations. Human evaluation can be time-consuming and resource-intensive, and the results can be susceptible to individual biases and expertise levels. To address these challenges, some researchers have explored employing advanced LLMs to replace human evaluation of medical LLMs. These advanced LLMs are trained on vast datasets and possess a knowledge base comparable to that of humans. Since the LLM evaluates the performance of medical LLMs from a human perspective, it is categorized under Human-Centric Evaluation.
4.2.2 Evaluation Dimensions
@@
In Human-Centric Evaluation, the initial step involves identifying the evaluation dimensions for the model. Given the inconsistent terminology employed in previous studies of medical LLMs to describe these dimensions, we standardized the terminology and provided corresponding explanations. Tab. 3 uses the normalized terms.
@@
Doctor-like: Exhibits a professional physician’s communication style and tone, reflecting a comprehensive understanding of medical principles and clinical expertise. Capable of providing clear explanations of intricate medical concepts and offering well-founded diagnostic insights and recommendations.
@@
Patient-friendly: Interacts with users in a patient and empathetic manner, fostering a sense of understanding and support. Effectively simplifies complex medical concepts into easily comprehensible language, demonstrating compassion and care, alleviating user anxiety and enhancing trust and satisfaction.
@@
Professional: Exhibits a high degree of professionalism, characterized by comprehensive medical knowledge and sound clinical judgment. Provides treatment recommendations and prescriptions that are not only accurate and reliable but also conform to medical norms and ethics, ensuring the achievement of optimal patient outcomes.
@@
Safe: Responses are safe, harmless, and not geographically or racially discriminatory. Correctly address unhealthy and unsafe issues, provide accurate and reliable advice, and avoid misleading users or raising health risks. Always follow medical ethics, protect user privacy, and ensure the accuracy and safety of information delivery.
@@
Helpful: Capable of meeting the user’s needs for medical consultation or the physician’s clinical instructions to achieve users’ expectations.
Given the critical importance of safety in the medical field, the safety dimension must be assessed by medical experts. Other dimensions can be evaluated by LLMs, thereby increasing efficiency. For non-professional users, the evaluation dimensions focus on patient-friendly and fluency. These two dimensions match the user’s perspective and can measure the real feedback of the model from users in real scenarios.
4.2.3 Protocols
Human-centric evaluation comprises two primary protocols: individual evaluation and pairwise evaluation.
Individual evaluation involves a human evaluator or an LLM scoring the output based on predefined dimensions, with only a single question and its corresponding output visible. The scoring range is a predetermined interval, and the average score for each dimension is statistically calculated as the final result. This metric quantitatively reflects the model’s overall performance in each dimension, identifying strengths and weaknesses and guiding future optimizations on weaker dimensions.
Pairwise evaluation involves a human evaluator or an LLM comparing the outputs of two models based on predefined dimensions and selecting the better-performing one. Evaluation metrics include win rate (the proportion of evaluated models outperforming the comparison model), tie rate (the proportion of the two that are equivalent), and loss rate (the proportion of evaluated models under-performing the comparison model). In addition, to minimize the positional bias existing in the LLM as a judge, the positions of the two model outputs are exchanged for another comparison. Pairwise evaluation facilitates the rapid identification of the best-performing model from multiple candidates, serving as a foundation for model optimization and deployment.
@@ 4.2.3 Protocols
5 Challenges and Future Directions
@@
5.1 Proactive Questioning Ability
Compared to LLMs in other domains, medical LLMs must not only provide accurate responses to medical inquiries but also possess the capability to proactively guide users in describing their concerns clearly. Users often struggle to convey their conditions due to a lack of medical expertise. Thus, medical LLMs should ask users about unclear information and help them articulate the underlying causes of their illnesses. Additionally, users may provide incorrect descriptions of the symptoms of the disease, necessitating that medical LLMs correct these misunderstandings. Therefore, the challenge is how to enhance the proactive questioning abilities of medical LLMs.
To address this, Bianque [23] constructed a multi-round conversation corpus for model training, which balances the abilities to answer and ask questions, thereby improving the proactive questioning capability of medical LLMs at the corpus level. Currently, researchers may consider exploring the design of agent workflows. For instance, the medical counseling task can be decomposed into three components: (1) assessing the level of detail in the user’s question, (2) identifying what information is missing, and (3) either answering the question or asking for more details from the user. Additionally, incorporating agent paradigms such as reflection and tool use into this workflow can enhance the credibility of responses generated by medical LLMs.
5.2 Multi-organization joint training
@@
To enable medical LLMs to achieve capabilities comparable to those of medical professionals, it is essential to integrate more real medical data into these models to enhance their medical knowledge. Data from professional medical organizations tend to be more authentic and reliable than data sourced from the web. However, due to the sensitive nature of medical data and associated policies, such data is seldom publicly available [49]. Consequently, a significant challenge is to ensure the secure utilization of medical data within organizations while safeguarding patient privacy. One approach to addressing this issue is federated learning [106], a distributed training method that enables model parameters to be shared across multiple organizations without sharing the data itself. By storing data locally, federated learning preserves data privacy while allowing the model to learn from diverse data sources, thus enhancing its generalization and performance. Federated learning has been applied to medical images [186], sensor data [187], and natural language text [117], promoting the secure use of health data among medical organizations.
@@
Currently, federated learning has several shortcomings in the medical field that require further investigation. Differential privacy [37] can be integrated with federated learning to enhance data privacy protection, but this integration degrades model performance. How to balance privacy preservation and model performance in the medical LLMs training is a worthy research topic. Additionally, the quality and quantity of data varies among healthcare organizations, and how to evaluate the contribution of each organization is a crucial topic. The establishment of a standardized set of evaluation metrics to quantify the impact of different data sources could be considered. Finally, medical LLMs have a significantly higher number of parameters than pretrained language models. Researchers may consider optimizing communication and computational overhead during the transmission and aggregation of parameters in large-scale models. For example, explore how to combine parameter-efficient fine-tuning methods with federated learning.
5.3 Personalized Diagnosis